OpenStudy (anonymous):

A certain bacteria has an exponential growth pattern. after 3 hours there are 51 bacteria, and after 12 hours there are 685 bacteria. how many bacteria will be present after 24 hours?

OpenStudy (lgbasallote):

when t = 3.. x =\(x_o\) = 51 then when t = 12...x = 685 therefore \[\ln(\frac{x}{x_o} )= kt\] \[\ln (\frac{685}{51} = k(12 -3)\] \[\ln (13.4313) = k(9)\] \[\frac{\ln (13.4313)}{9} = k\] \[k = 0.2886\] now use \[x = x_oe^kt\] \[x = 51 e^{(0.2886)(24-3)})\] \[x = 51e^{6.0606}\] \[x = 51(428.6325)\] \[x = 21,860.25\] \[x = 21,860\] got it?